OpenStudy (anonymous):

The nucleus of a 125Xe atom (an isotope of the element xenon with mass 125 u) is 6.0 fm in diameter. It has 54 protons and charge q =+54e. What is the electric force on a proton 1.8fm from the surface of the nucleus? Hint: Treat the spherical nucleus as a point charge. Express your answer to two significant figures and include the appropriate units. What is the proton's acceleration? Express your answer to two significant figures and include the appropriate units.

OpenStudy (anonymous):

@ganeshie8 or @Kainui.

OpenStudy (anonymous):

i found this solution online..... 9 × 10^9*(54) (1.602 × 10^−19)^2/(5 × 10^−15)^2, but not sure where the 1.602*10^-19 and 5*10^-15 came from

OpenStudy (kainui):

Show me your best attempt at solving it, and explain as much as you can, even if you think it's wrong so I can help you where you need it and let you know what you have right.

OpenStudy (anonymous):

well that's where i'm stuck. i get that q1 is 54 coulombs, but i have no idea where the q2 comes from and what the distance is. or what fm is.

OpenStudy (anonymous):

distance is gonna be the radius of the nucleus (6/2) plus the distance of the proton which is 4.8 fm, but i've never worked with that unit so i don't know what it is. or where they got 1.602*10^-19 for the second charge.

OpenStudy (anonymous):

after that i can plug and chug.

OpenStudy (kainui):

One question, what is the question asking for, the force of what on what? Also, fm is a femtometer (really tiny) just like cm, mm, and km are centimeter, milimeter, and kilometer.

OpenStudy (anonymous):

it wan'ts the force on the proton and then the acceleration. the second part shouldn't be an issue once i find force. F=ma.

OpenStudy (kainui):

Ok, so do you see how to find what "q2" and the distance should be now?

OpenStudy (anonymous):

not q2. but distance, yes. q1 is 54 coulombs, but i can't figure out q2. it's the charge of the proton, but how do i find that?

OpenStudy (anonymous):

or is that a constant value i'm supposed to just know?

OpenStudy (kainui):

Yeah, all protons are the same charge, in fact, the charge of a proton is exactly the opposite of an electron charge. So once you have it memorized you really have both memorized.

OpenStudy (anonymous):

oh, i don't have that memorized. lol. only second week of class. now, when i convert the mass from u, do i want g or kg? can't remember

OpenStudy (anonymous):

for the second part

OpenStudy (kainui):

usually you'll want it to be kg, but it depends, just make sure the units match up. A useful thing to know is \[\Large \left[ kg \frac{m}{s^2} \right] = [N]\]

OpenStudy (anonymous):

force =3.38*10^21=125u*a. solve for a.

OpenStudy (anonymous):

ok, thanks.

OpenStudy (anonymous):

that force seems large....

OpenStudy (anonymous):

is saying the charge of q1 +54e not just 54 coulombs?

OpenStudy (anonymous):

ok, nvmd. i got it. thanks for your help!